### v2

# DEIO: Deep Event Inertial Odometry

Abstract
Event cameras are bio-inspired, motion-activated sensors that demonstrate impressive potential in handling challenging situations, such as motion blur and high-dynamic range. Despite their promise, existing event-based simultaneous localization and mapping (SLAM) approaches exhibit limited performance in real-world applications. On the other hand, state-of-the-art SLAM approaches that incorporate deep neural networks for better robustness and applicability. However, these is a lack of research in fusing learning-based event SLAM methods with IMU, which could be indispensable to push the event-based SLAM to large-scale, low-texture or complex scenarios. In this paper, we propose DEIO, the first monocular deep event-inertial odometry framework that combines learning-based method with traditional nonlinear graph-based optimization. Specifically, we tightly integrate a trainable event-based differentiable bundle adjustment (e-DBA) with the IMU pre-integration in a factor graph which employs keyframe-based sliding window optimization. Numerical Experiments in nine public challenge datasets show that our method can achieve superior performance compared with the image-based and event-based benchmarks. The source code is available at: https://github.com/arclab-hku/DEIO.
Index Terms:
Event-based Vision, 6-DoF Pose Tracking, Event Camera, SLAM, Deep Learning, Robotics.
I INTRODUCTION
I-A Motivation
Achieving reliable and accurate Simultaneous Localization and Mapping (SLAM) under adverse conditions remains challenging, particularly when employing image-based solutions (RGB or RGB-D cameras). Event camera is a motion-activated sensor that only captures pixel-wise intensity changes and reports them as asynchronous stream instead of the whole scene at intensity image as a frame. Because of their remarkable properties, such as high temporal resolutions, high dynamic range (HDR), and no motion blur, event cameras have the potential to enable high-quality perception in extreme lighting conditions and high-speed motion scenarios that are currently not accessible to standard cameras. Nevertheless, integrating event cameras into SLAM systems presents significant challenges. This is caused by the fact that the event data is sparse, irregular, and asynchronous, with each event conveying limited information and containing inherent noise. What’s more, unlike standard-vision counterparts, event camera cannot capture visual information when motion is parallel to edges or in static state. Therefore, current purely event-based SLAM systems [1, 2] generally lack the robustness requirement for many real-world applications.
Using additional sensors tends to achieve superior performance since different modalities can complement each other, such as image frame [3, 4, 5], depth sensors [6], or even LiDAR [7]. However, these combination might limit the application of event cameras in real-world devices due to increased costs and more complicated sensor calibration requirements [8]. What’s more, they might introduce some bottlenecks, for example, relying the image frames might makes the system susceptible to motion blur and reduce the benefits of HDR.
Recent studies have introduced learning-based approaches [9] as promising solutions for event-based SLAM. These methods, which are solely trained end-to-end on synthetic events, demonstrate remarkable generalization to real-world benchmarks and can outperform some traditional systems in accuracy. However, it’s important to note that visual-only systems have inherent limitations, making them vulnerable to low-textured environments and suffering from scale ambiguity. To address these visual degradation, a promising strategy is to incorporate multiple sensors, the Inertial Measurement Unit (IMU), which is low-cost and readily available in most event cameras. What’s more, in the minimal configuration of a monocular setup, the IMU can also be used to recover the metric scale and provide higher accurate and more robust system. Despite this promising, learning-based event SLAM with IMU remains an unexplored territory.
I-B Contributions
Therefore, in this paper, we propose a deep learning-based event-inertial odometry (DEIO), which combines learning-based methods with traditional nonlinear graph-based optimization. We leverage the strengths of neural networks in predicting event correspondence, replacing the optical flow or hand-craft event-corner feature tracking [10]. To the best of our knowledge, DEIO is the first learning-based event-inertial odometry framework. More specifically, we design the learning-optimization-combined framework that tightly-coupled integrate trainable event-based differentiable bundle adjustment (e-DBA) with IMU pre-integration. In the framework, event-based recurrent optical flow and e-DBA are performed among consecutive event streams. The Hessian information is derived from e-DBA and further fed into a generic factor graph, facilitating integration with the IMU through keyframe-based sliding window optimization. The e-DBA is trained on synthetic event data but demonstrate strong generalization on real-world challenge scenarios. Our main contributions can be summarized as follows:
  1. 1.
    We propose a learning-based event-inertial odometry, which integrates a trainable e-DBA and IMU information through graph-based optimization, which improves the accuracy and robustness of pose estimation.
  2. 2.
    DEIO is the first learning-based monocular event-inertial odometry. Even though training on synthetic data, it outperforms over 20 state-of-the-art methods across 9 challenging real-world event benchmarks. The performance gap is further increased when trained on real data.
  3. 3.
    We release our code and the preprocessed event data to facilitate further research in learning-based event pose tracking.
II Related Work
II-A Traditional Approaches for Event-based SLAM
The first purely event-based VO was presented in [11], which performed real-time event-based SLAM through three decoupled probabilistic filters that jointly estimate the 6-DoF camera pose, 3D map of the scene, and image intensity. However, it is computationally expensive and requires GPU to achieve real-time performance. EVO [1] presents a geometric 2D-3D model alignment, utilizing the 2D image-like event representation and the 3D map reconstructed from EMVS[12]. It can perform on small-scale scenes, but it is highly parameter-sensitive and requires to run in the scene that is planar to the sensor, up to several seconds, for bootstrapping the system. ESVO [2] is the first stereo event-based VO method to estimate the ego motion through the 2D-3D edge registration on time surface. However, it barely operates in real-time in low resolution event camera and also faces limitations due to rigorous initialization as well as unreliable pose tracking. What’s more, some approaches reply on a pre-build photometric 3D map [13] or restrict the motion type to rotation-only [14]. However, most of them are prone to failure in challenging scenarios (such as fast motion), which negates the unique advantages of event cameras.
To enhance the robustness of purely event-based VO, existing event-based SLAM methods have demonstrated good performance by incorporating additional sensors. Notably, event-inertial integration is a widely used approach to address the limitations of event-only SLAM, which provides scale awareness and continuity of estimation with minimal setup requirements. A monocular event-inertial odometry (EIO) is presented in [10], which employs the event-corner features with IMU measurement to deliver real-time and accurate 6-DoF state estimation. What’s more, this EIO framework can provide good global consistency through loop closure of the event-based representation. Ultimate-SLAM [3] and PL-EVIO [4] investigate the complementary nature of events and images to present an event-image-IMU odometry (EVIO). Chamorro et al. [15] integrate the event-based line feature with the IMU to achieve high ultra-fast pose estimation. ESVIO [16] proposed the first stereo EIO and EVIO framework to estimate states through temporally and spatially event-corner feature association. EVI-SAM [17] introduced the first event-based hybrid pose tracking framework, merging the feature-based and direct-based methods. Additionally, it consistently delivers excellent mapping results in a variety of challenging scenarios, demonstrating performance comparable to learning-based and NeRF-based mapping methods. Despite such remarkable success, current event-based SLAM systems still fall short of the robustness required for certain real-world applications, and their performance has hit a bottleneck, this directs toward the exploration of deep learning approaches.
II-B Learning-based Approaches for Event-based SLAM
Zhu et al. [18] constructed a convolutional neural networks (CNN) with an unsupervised encoder-decoder architecture for semi-dense depth prediction. Hidalgo-Carrió et al. [19] was the first to estimate dense depth maps from a monocular event camera through recurrent CNN architecture. Gehrig et al. [20] apply recurrent neural networks (RNN) architecture which maintained an internal state that was updated through asynchronous events or irregular images and could be queried to dense depth estimation at any timestamp. Ahmed et al.[21] developed a deep event stereo network that reconstructs spatial image features from embedded event data and leverages the event features using the reconstructed image features to compute dense disparity maps. However, all of these methods require knowledge of camera motion and show poor generalization beyond the training scenarios.
DH-PTAM [22] proposed a stereo event-image parallel tracking and mapping system that replacing hand-crafted features with learned features. Klenk, et al. [9] extends the DPVO [23] to accommodate the event modality through voxel-based representation, which are able to demonstrate generalization from simulation to several real-world event-based benchmarks. It shows good performance compared to both classical and deep-learning baselines. However, their results reveal a significant scale loss/ambiguity, and they requires the careful selection of the input data. Pellerito, et al. [24] introduces an end-to-end VO which also builds upon DPVO using complicated feature encoders to fuse asynchronous events with image data. However, it requires a minimum of 1.2 million events between each pair of frames and discards additional frames if this requirement is not met. More importantly, it failed to evaluate their methods under fast motion or HDR scenes, and could only operate in very limited amount of real event data without challenging situations.
Our framework is closely related to [23, 9] but extends these works by integrating trainable event representations with IMU through our learning-optimization-combined strategy to improve the performance and robustness. Additionally, it is important to note that a visual-only (event-only) system, such as DPVO [23] and DEVO [9] has inherent limitations, failures can manifest in numerous ways, such as in low-texture environment, divergence in the optimization algorithm, scale ambiguity, and etc. Nevertheless, there is a absence of an implementation that incorporates learning-based event SLAM with IMU fusion framework.
III Methodology
The overall design of the framework aims to tightly-couple fuse the trainable eDBA with IMU pre-integration through optimization in a keyframe-based sliding window. Fig. 1 depicts the overview of our system. The main innovations of this work reside in how learning-based methods and traditional graph optimization are fused together. For the front-end, a deep neural network (Section III-A) is utilized to compute sparse patch-based correspondence for event-based recurrent optical flow. On the back-end, the re-projections are used to formulate an e-DBA problem, while the Hessian information derived from the e-DBA (detailed in Section III-C) is then tightly fused with the IMU pre-integration through a patch-based co-visibility factor graph (Section III-D).
Figure 1: Overview of the DEIO system. Our system builds upon the event-based recurrent optical flow network from [23, 9], and introduces a learning-optimization combined framework to enhance robustness and accuracy of state estimation. It utilizes a patch-based co-visibility factor graph to tightly-coupled fuse the Hessian information from the e-DBA and the IMU pre-integration using event-IMU combined bundle adjustment.
III-A Network Structure for Event-based Optical Flow
In our framework, the event-based optical flow leverages the patch graph and recurrent network architecture from DPVO [23], which is a sparse patch-based VO derived from well-developed optical flow network (RAFT [25]). Additionally, we incorporate the voxel-based event representation from [9] to effectively integrate the event modality into the DPVO. Each event voxel grid represents a frame-like structure of the event stream. The structure of the event-based recurrent network is designed to estimate the 2D motion of the event patches across time by predicting optical flow. While the estimated optical flow is integrated into the differentiable bundle adjustment layer to update the depth of the event-patches and the camera poses. The structure for estimating event-patch trajectory consists of three primary components: (i) The CNN network that extracts patch-based event feature representations, capturing local context through matching and contextual features; (ii) The correlation layer that computes visual similarity between these event patches; (iii) The recurrent update operator that handles these event-patches correspondence coupled with differentiable bundle adjustment. These construct the e-DBA problem, enabling end-to-end learning of reliable event-based optical flow, and optimize pose as well as depth of each patch iteratively through a differentiable bundle adjustment layer by minimizing the re-projection residual. The learning phase of Fig. 1 illustrates the complete process.
By dividing the event stream between frame $i-1$ and $i$ into $d$ equally spaced time intervals and aggregating each interval, an event voxel grid $\mathbf{E}_{i}\in\mathbb{R}^{d\times H\times W}$ is obtained. After that, sparse event-patches are extracted to build the patch graph for scene representation. We denote the event patches for frame $i$ contains a series of $\bm{P}_{in}$ with size of $p\times p$:
$\bm{P}_{in}=\left[\begin{matrix}\bm{x}_{i}\\ \bm{y}_{i}\\ 1\\ \bm{d}_{i}\\ \end{matrix}\right]\quad\bm{x}_{i},\bm{y}_{i},\bm{d}_{i}\in\mathbb{R}^{1\times p^{2}}$ (1)
where $\bm{d}_{i}$ is the estimated inverse depth, assuming a constant depth for each individual event patch. The depth of each event patch is initialized using the median depth of all patches extracted from the previous three frames and would be further optimized in the e-DBA layer. We denote the $n^{th}$ event patches for frame $i$ as $\bm{P}_{in}$ with totally patch number $N$ ($N=96$ in our implementations). Given the camera pose $\bm{T}_{i}$ in frame $i$, the camera intrinsics $\bm{\pi}$, the event patch can be re-projected to other frame $j$ and reconstruct the edges from $\bm{P}_{in}$ as $(n,j)$, the edge-based patch graph $\mathcal{G}$. Thus the optimized pose $\bm{T}_{ij}$ and depth $\bm{d}_{i}$ can be obtain by minimizing the following cost function:
$e=\sum_{i}\sum_{(n,j)\in\mathcal{G}}||\bm{F}(\bm{T}_{ij},\bm{d}_{i})||^{2}_{\bm{\Sigma}_{inj}}$ (2)
$=\sum_{i}\sum_{(n,j)\in\mathcal{G}}||\bm{\pi}[\bm{T}_{j}^{-1}\cdot\bm{T}_{i}\cdot\bm{\pi}^{-1}(\hat{\bm{P}}_{in})]-[\hat{\bm{P}_{in}}+\bm{\delta}_{inj}]||^{2}_{\bm{\Sigma}_{inj}}$
where $\bm{\pi}$ and $\bm{\pi}^{-1}$ is the projection and back-projection function of the event camera. $(\bm{\delta}_{inj},\bm{\Sigma}_{inj})$ is the patch-based optical flow field predicted by the event-based recurrent network. The term $\bm{\delta}_{inj}\in\mathbb{R}^{2}$ denotes a 2D flow vector that indicates how the re-projection of the event-patch center should be updated, while the $\bm{\Sigma}_{inj}$ serves as the patch-wise confidence weight of the optical flow. $\hat{\mathbf{P}}_{in}$ denotes the center of the patch $\mathbf{P}_{in}$. $\bm{T}_{i}$ and $\bm{T}_{j}$ is the event camera pose in $\mathbb{SE}(3)$. Outlier rejection is handled automatically by the network, enhancing robustness compared to classical event-based SLAM systems.
III-B Training and Supervision for the e-DBA
The learning process for e-DBA of DEIO leverages supervision from poses, flow, and patch selection scores. The pose loss, $\mathcal{L}_{\text{pose}}$, is calculated as follows:
$\mathcal{L}_{\text{pose}}=\sum_{i=1}^{N}\left\|\log_{\mathbb{SE}(3)}(T_{\text{GT},ij}^{-1}\bm{T}^{{}^{\prime}}_{ij})\right\|$ (3)
where $\bm{T}_{\text{GT},ij}$ and $\bm{T}^{{}^{\prime}}_{ij}$ are the ground truth and updated poses of e-DBA between timestamp $i$ and $j$ respectively ($\bm{T}^{{}^{\prime}}_{ij}=\bm{T}_{i}^{-1}\bm{T}_{j}^{{}^{\prime}}$). The flow loss, $\mathcal{L}_{\text{flow}}$, measures the average distance between the estimated and ground truth optical flow over edges in the patch graph, where the frame distance between the source and target frames of each event-patch is within two frames. Regarding the event-patch selection, given the sparsity of event data in voxel, to avoid selecting event-patch from event-free regions, we follow [9] by adopting a patch selection network to propose regions for patch extraction and the loss on selection scores $\mathcal{L}_{\text{score}}$. The overall loss is then defined as a weighted combination: $\mathcal{L}=10\mathcal{L}_{\text{pose}}+0.1\mathcal{L}_{\text{flow}}+0.05\mathcal{L}_{\text{score}}$.
The e-DBA is implemented in PyTorch and trained for 240K steps on two NVIDIA RTX-3090 GPUs. Each step uses a batch size of 1 with a sequence of 15 frames where the first 8 frames are used for initialization, and the remaining 7 frames are incrementally added for training. During training, each frame contains 80 event patches, each sized 12px×$\times$12px. In the first 1000 training steps, the network estimates only the depth of patches with fixed poses, allowing it to stabilize early training. In the subsequent steps, it jointly estimates both pose and patch depth. In the edge state updater, we conduct 18 iterations, supervising the pose and flow at each iteration, while the selection score is supervised only in the final iteration. Regarding the optimization objective Eq. 2 in e-DBA, we perform two Gauss-Newton iterations to refine poses and patch depths, ensuring that the predicted flow aligns with the camera motion.
Since the recurrent optical flow pipeline has demonstrated strong generalization [25, 26, 23, 9], we utilize the network weights pre-trained on the synthetic event-based TartanAir [27] dataset using the ESIM [28] simulator for all subsequent tests in this work. To further bridge the sim-to-real gap, we fine-tune the pre-trained model by incorporating real data into the training process. The resulting performance improvements would be further discussed in Section IV-D
III-C Hessian Derived from the e-DBA
The linearized expression of the re-projection errors in Eq. (2) can be written as:
$\delta\bm{F}(\bm{T}_{ij},\bm{d}_{i})=\left[\begin{matrix}\bm{J}_{ij}&\bm{J}_{\bm{d}_{i}}\end{matrix}\right]\left[\begin{matrix}\bm{\xi_{ij}}\\ \Delta\bm{d}_{i}\end{matrix}\right]$ (4)
where $\delta\bm{F}(\bm{T}_{ij},\bm{d}_{i})\in\mathbb{R}^{2}$ is the residual matrix with size of $2NK\times 1$. The $\bm{\xi_{ij}}$ is the Lie algebras of the updated pose $\bm{T}_{ij}$ in $\mathbbm{se}(3)$. $\Delta\bm{d}_{i}$ denotes the error state of the inverse depth. The Jacobians $\bm{J}_{ij},\bm{J}_{\bm{d}_{i}}$ are with respect to the update pose $\bm{T}_{ij}$ and depth $\Delta\bm{d}_{i}$, respectively. An event-patch $\bm{P}_{in}$ can be re-projected from frame $i$ into frame $j$ follows the warping function:
$\bm{P}^{{}^{\prime}}_{jn}=\left[\begin{matrix}x_{j}&y_{j}&z_{j}&w_{j}\end{matrix}\right]^{T}=\bm{\pi}[\bm{T}_{j}^{-1}\cdot\bm{T}_{i}\cdot\bm{\pi}^{-1}(\hat{\bm{P}}_{in})]$ (5)
Thereofre, the Jacobians of the re-projection can be expressed as:
$\bm{J}_{ij}=\left[\begin{matrix}\frac{f_{x}w_{j}}{z_{j}}&0&-\frac{f_{x}x_{j}w_{j}}{z_{j}^{2}}&-\frac{f_{x}x_{j}y_{j}}{z_{j}^{2}}&f_{x}+\frac{f_{x}x_{j}^{2}}{z_{j}^{2}}&-\frac{f_{x}y_{j}}{z_{j}^{2}}\\ 0&\frac{f_{y}w_{j}}{z_{j}}&-\frac{f_{y}y_{j}w_{j}}{z_{j}^{2}}&-f_{y}-\frac{f_{y}y_{j}^{2}}{z_{j}^{2}}&\frac{f_{y}x_{j}y_{j}}{z_{j}^{2}}&-\frac{f_{y}x_{j}}{z_{j}^{2}}\end{matrix}\right]$
where $f_{x}$ and $f_{y}$ are the given event camera intrinsic parameters. While the $\bm{J}_{\bm{d}_{i}}$ can be given as:
$\bm{J}_{\bm{d}_{i}}=\left[\begin{matrix}f_{x}(\frac{\bm{t}_{ij}[0]}{z_{j}}-\frac{\bm{t}_{ij}[2]x_{j}}{z_{j}^{2}})\\ f_{y}(\frac{\bm{t}_{ij}[1]}{z_{j}}-\frac{\bm{t}_{ij}[2]y_{j}}{z_{j}^{2}})\end{matrix}\right]$ (7)
where the $\bm{t}_{ij}$ is the translation vector of the related transform between $\bm{T}_{i}$ and $\bm{T}_{j}$. Then, the Hessian Matrix of Eq. (2) can be computed following:
$\bm{H}=\left[\begin{matrix}\bm{J}_{ij}&\bm{J}_{\bm{d}_{i}}\end{matrix}\right]^{T}\bm{W}\left[\begin{matrix}\bm{J}_{ij}&\bm{J}_{\bm{d}_{i}}\end{matrix}\right]\quad\bm{W}=diag(\bm{\Sigma}_{inj})$ (8)
By decoupling the pose and depth variables, the system can be efficiently solved using the Schur complement:
$\bm{H}\left[\begin{matrix}\bm{\xi_{ij}}\\ \Delta\bm{d}_{i}\end{matrix}\right]=-\left[\begin{matrix}\bm{J}_{ij}&\bm{J}_{\bm{d}_{i}}\end{matrix}\right]^{T}\bm{W}\bm{F}(\bm{T}_{ij},\bm{d}_{i})$ (9)
$\left[\begin{matrix}\bm{B}&\bm{E}\\ \bm{E}^{T}&\bm{C}\end{matrix}\right]\left[\begin{matrix}\bm{\xi_{ij}}\\ \Delta\bm{d}_{i}\end{matrix}\right]=\left[\begin{matrix}\bm{v}\\ \bm{u}\end{matrix}\right]$
and leading to the following equation:
$\bm{\xi_{ij}}=\underbrace{[\bm{B}-\bm{E}\bm{C}^{-1}\bm{E}^{T}]^{-1}}_{\bm{Hgg}}\underbrace{(\bm{v}-\bm{E}\bm{C}^{-1}\bm{u})}_{\bm{Vgg}}$ (10)
where $\bm{v}$ is a $6\cdot K\times 1$ vector (with 6 DoF pose and totally $K=10$ keyframe), and $\bm{u}$ is an $N\cdot K\times 1$ vector (with totally number of $N=96$ event patches for each frame). The $\bm{B}$ is the matrix with size of $60\times 60$, $\bm{E}$ is the residual matrix with size of $60\times 960$, and $\bm{C}$ is a diagonal matrix with size of $960\times 960$. A damping factor of $10^{-4}$ is also applied to $\bm{C}$ as [23]. These matrices establish an inter-frame pose constraint (represented by $\bm{Hgg}$ and $\bm{Vgg}$) that integrates the e-DBA information. After updating the camera poses $\bm{T}_{j}^{{}^{\prime}}=\bm{T}_{i}\cdot\bm{T}^{{}^{\prime}}_{ij}$, $\bm{T}_{ij}=exp(\bm{\xi_{ij}})$, the inverse depth of each event patch would be updated as:
$\bm{d}_{i}^{{}^{\prime}}=\Delta\bm{d}_{i}+\bm{d}_{i}$ (11)
$\Delta\bm{d}_{i}=\bm{C}^{-1}(\bm{u}-\bm{E}^{T}\bm{\xi_{ij}})$
The calculations of Eq. (10) and Eq. (11) can be efficiently performed in parallel on GPU with CUDA acceleration. While all the Hessian information $\bm{Hgg}$ and the corresponding $\bm{Vgg}$, derived from the co-visibility graph, are integrated into the factor graph where they are optimized on the CPU. The e-DBA contributes extensive geometric information, incorporating learned uncertainties, to the factor graph. While the optimization results (updated poses and depths) are then iteratively fed back to refine/re-weight the event-based optical flow network. This recurrent process enhances the robustness and accuracy of the overall system.
III-D Graph-based e-DBA and IMU Optimization
Unlike end-to-end approaches that use networks to fuse the features from two modalities (event and IMU) and predict poses directly, our DEIO combine the neural networks with event-inertial bundle adjustment. To this end, we design a learning-optimization-combined framework that tightly-coupled integrate the Hessian information from e-DBA and IMU pre-integration within keyframe-based sliding window optimization. The full state vector in the sliding window, with $k$ ( total number of keyframes $K=10$ in our implementation), is defined as:
$\bm{\chi}=[\bm{T}^{w}_{b_{k}},\bm{v}^{w}_{b_{k}},\bm{b}_{a_{k}},\bm{b}_{g_{k}}],k=1,2,3,...$ (12)
where $\bm{T}^{w}_{b_{k}}=\left[\begin{matrix}\bm{R}^{w}_{b_{k}}&\bm{t}^{w}_{b_{k}}\\ 0&1\end{matrix}\right]\in\mathbb{SE}(3)$ is the pose of the body (IMU) frame in the world frame, given by the translation $\bm{t}^{w}_{b_{k}}$ and rotation matrix $\bm{R}^{w}_{b_{k}}$; $\bm{v}^{w}_{b_{k}}$ is the velocity of the IMU in the world frame; $\bm{b}_{a_{k}}$ and $\bm{b}_{g_{k}}$ is the accelerometer bias and gyroscope bias, respectively.
We solve the state estimation problem by constructing a factor graph with the GTSAM library and optimizing it with the Levenberg-Marquardt. The cost function, which minimizes residuals from various factors corresponding to different aspects of data constraints, can be written as:
$\bm{J}(\bm{\chi})=||\bm{r}^{k}_{event}||^{2}_{W^{k}_{event}}+\sum_{k=0}^{K-1}||\bm{r}^{k}_{imu}||^{2}_{W^{k}_{imu}}+||\bm{r}_{m}||^{2}_{W_{m}}$ (13)
Eq.(13) contains the event-based residuals $\bm{r}^{k}_{event}$ with weight $W^{k}_{event}$; the IMU pre-integration residuals $\bm{r}^{k}_{imu}$ with weight $W^{k}_{imu}$; the marginalization residuals $\bm{r}_{m}$ with weight $W_{m}$;
Given the Hessian information $\bm{Hgg}$ and the corresponding $\bm{Vgg}$, the event residual factor can be written like [29]:
$\bm{r}^{k}_{event}=$ $\frac{1}{2}\left[\begin{matrix}\bm{\xi^{w}_{e_{0}}}&\bm{\xi^{w}_{e_{1}}}&\cdots&\bm{\xi^{w}_{e_{k}}}\end{matrix}\right]\bm{Hgg}\left[\begin{matrix}\bm{\xi^{w}_{e_{0}}}\\ \bm{\xi^{w}_{e_{1}}}\\ \vdots\\ \bm{\xi^{w}_{e_{k}}}\end{matrix}\right]$ (14)
$-\left[\begin{matrix}\bm{\xi^{w}_{e_{0}}}&\bm{\xi^{w}_{e_{1}}}&\cdots&\bm{\xi^{w}_{e_{k}}}\end{matrix}\right]\bm{Vgg}$
where $\bm{\xi}^{w}_{e_{k}}=\bm{\xi}^{e}_{b}\cdot\bm{\xi}^{w}_{b_{k}}$, and $\bm{\xi}^{w}_{b_{k}}=\log_{\mathbb{SE}(3)}(\bm{T}^{w}_{b_{k}})$. While $\bm{\xi}^{w}_{e_{k}}$ and $\bm{\xi}^{w}_{b_{k}}$ are the Lie algebras of the event camera pose and IMU pose in $k^{th}$ keyframe, respectively. $\bm{\xi}^{e}_{b}$ is the extrinsics of the event camera and IMU.
Eventually, the IMU residual factor can be derived as follows:
$\begin{split}\bm{r}^{k}_{imu}=\left[\begin{array}[]{c}\bm{R}^{b_{k}}_{w}(\bm{t}^{w}_{b_{k+1}}-\bm{t}^{w}_{b_{k}}-\bm{v}^{w}_{b_{k}}\Delta t-\frac{1}{2}\bm{g}^{w}\Delta t^{2})-\hat{\bm{\alpha}}^{b_{k}}_{b_{k+1}}\\ \bm{R}^{b_{k}}_{w}(\bm{v}^{w}_{b_{k+1}}-\bm{v}^{w}_{b_{k}}-\bm{g}^{w}\Delta t)-\hat{\bm{\beta}}^{b_{k}}_{b_{k+1}}\\ 2\left[(\bm{q}^{w}_{b_{k}})^{-1}\otimes\bm{q}^{w}_{b_{k+1}}\otimes(\hat{\bm{\gamma}}^{b_{k}}_{b_{k+1}})^{-1}\right]_{xyz}\\ \bm{b}_{a_{k+1}}-\bm{b}_{a_{k}}\\ \bm{b}_{g_{k+1}}-\bm{b}_{g_{k}}\\ \end{array}\right]\end{split}$ (15)
where $\bm{\alpha}^{b_{k}}_{b_{k+1}}$, $\bm{\beta}^{b_{k}}_{b_{k+1}}$, $\bm{\gamma}^{b_{k}}_{b_{k+1}}$ are the IMU pre-integration term [4]; $\bm{g}^{w}$ is the gravity vector; $\Delta t$ is the time interval between keyframe $k$ and $k+1$; $\bm{q}^{b_{k}}_{w}$ is the quaternion of the corresponding rotation matrix $\bm{R}^{b_{k}}_{w}$ with $[\cdot]_{xyz}$ extracts the vector portion.
Figure 2: Patch-based co-visibility factor graph
III-E EIO System
As illustrated in Fig. 2, our DEIO maintains a patch-based co-visibility factor graph that takes the raw event stream and IMU data as input, and perform 6 DoF pose estimation.
Initialization
We use 8 frames for event-only initialization. The new event patches are added until 8 frames are accumulated, and then run 12 iterations of the update operator, followed by two event-only bundle adjustment iterations. Since the IMU provides scale awareness and the gravity direction, it is essential to initialize the states in a metric-scale local frame for better convergence. Similar to the initialization procedure from [30, 31], where the vision-only structure from motion (SfM) is replaced by the event-only bundle adjustment, after the event-only initialization, we further employ the IMU pre-integration to establish the up-to-scale camera pose. After linear alignment and gravity refinement, the scale between IMU and event camera is restored, with further correction applied to the depth of the event-based patch. Subsequently, the event-IMU bundle adjustment (Section III-D) is iteratively performed by updating the co-visibility graph based on the event-inertial initialization results. To enable initialization, enough camera motion is required. During the event-only initialization, we accumulate only frames that exhibit an average optical flow magnitude of at least 8 pixels (estimated by performing one update iteration) from the prior frame. What’s more, sufficient excitation is also required from the IMU, with the gyroscope’s variance exceeding 0.25.
Co-visibility Patch Graph Management
When a new event patch is added, the edges between that event patch and the previous $r$ keyframes ($r=13$ in our implementations) are added into the patch graph. After appending these new edges (with the predicted event-based optical flow field $(\bm{\delta}_{inj},\bm{\Sigma}_{inj})$), the update operation of the patch graph is performed, including recurrent update operator and the 2 event-IMU bundle adjustment iterations. After that, the keyframe selection is perform to maintain the keyframe-based sliding window.
Keyframe Strategy
The most recent 3 frames are always considered as keyframes. After each update operation of the co-visible graph, the optical flow magnitude between the $(t-5)^{th}$ and $(t-3)^{th}$ keyframe would be calculated. If this magnitude is less than 60 pixels, the $(t-4)^{th}$ keyframe, along with all associated edges and patch-based features, would be removed. Additionally, the event patches and keyframe are discarded once they fall outside the optimization window. The oldest keyframe that is abandoned would be marginalized, along with its associated IMU pre-integration, to construct the marginalization factor $\bm{r}_{m}$ in Eq. (13).
IV Experiments
We conducted quantitative and qualitative evaluations of our DEIO across nine challenging real-world datasets with varying camera resolution and diversity scenarios on different platforms. Specifically, in subsections IV-B and IV-C, we assess the performance of our DEIO in night driving scenarios and indoor low-light quadrotor flights, respectively. DEIO is compared against over 20 baseline methods from various literature, including both event-based and image-based approaches. Additionally, we proposed three additional baselines: (i) DPVO [23], which operates on the RGB frames of the respective dataset; (ii) DBA-Fusion [29], a learning-based VIO framework that combines DROID-SLAM [26] with IMU data, which runs on the RGB frames and IMU; (iii) a modified version of our proposed learning-optimization-combined framework that uses images and IMU data as inputs, referred to as DVIO. It operates on the standard images and IMU data from the respective datasets, with the network weights provided by [23]. This enables a comparison between events and images within the same framework and parameters.
To ensure fair comparison, same trajectory alignment protocol is required, therefore, we had to employ different alignment ways and evaluation criterion according to the compared baseline methods. For example, regarding the DAVIS240c dataset [32], the estimated and ground-truth trajectories were aligned using a 6-DOF transformation (in SE3) over a 5-second interval [0-5s], and we calculated the mean position error (MPE) as a percentage of the total traveled distance of the ground truth as the evaluation criterion. In contrast, for the stereo HKU dataset [16], although MPE is also used as the metric, to maintain consistency with the baseline methods, we aligned the estimated trajectories and ground truth in SE3 using all available data rather than just the 5-second segment, While for the TUM-VIE dataset [33], the baseline methods calculated the root-mean-square (RMSE) of the Absolute Trajectory Errors (ATE) based on the estimated camera pose as the evaluation criterion. To avoid confusion, we would annotate the trajectory alignment protocol in the footnotes of each table, ensuring that different methods within the same table utilize the same alignment protocol. What’s more, the notations E, F, and I in each table represent the use of event, frame, and IMU, respectively, which helps facilitate a better performance comparison among different modalities.
IV-A Comparison with SOTA Methods in Challenge Benchmarks
We only choose the challenge enough real-world event dataset for our evaluations, and the details of these benchmark evaluation are as follows:
TABLE I: Accuracy Comparison of Our DEIO with Other Monocular Event-based Baseline in DAVIS240c Dataset [32]
Methods Modality boxes_translation hdr_boxes boxes_6dof dynamic_translation dynamic_6dof poster_translation hdr_poster poster_6dof Average
Zhu et al. [34] E+I 2.69 1.23 3.61 1.90 4.07 0.94 2.63 3.56 2.58
Henri et al. [35] E+I 0.57 0.92 0.69 0.47 0.54 0.89 0.59 0.82 0.69
Ultimate-SLAM [3] E+I 0.76 0.67 0.44 0.59 0.38 0.15 0.49 0.30 0.47
E+F+I 0.27 0.37 0.30 0.18 0.19 0.12 0.31 0.28 0.25
Jung et al. [36] E+I 1.50 2.45 2.88 4.92 6.23 3.43 2.38 2.53 3.92
E+F+I 1.24 1.15 0.98 0.89 0.98 1.83 0.57 0.97 1.07
HASTE-VIO [37] E+I 2.55 1.75 2.03 1.32 0.52 1.34 0.57 1.50 1.45
EKLT-VIO [38] E+F+I 0.48 0.46 0.84 0.40 0.79 0.35 0.65 0.35 0.54
Dai et al. [39] E+I 1.0 1.8 1.5 0.9 1.5 1.9 2.8 1.2 1.56
Mono-EIO [10] E+I 0.34 0.40 0.61 0.26 0.43 0.40 0.40 0.26 0.39
PL-EVIO [4] E+F+I 0.06 0.10 0.21 0.24 0.48 0.54 0.12 0.14 0.24
Lee et al. [40] E+F+I 0.74 0.69 0.77 0.71 0.86 0.28 0.52 0.59 0.65
EVI-SAM [17] E+F+I 0.11 0.13 0.16 0.30 0.27 0.34 0.15 0.24 0.21
DPVO [23] F 0.02 0.71 0.59 0.09 0.05 0.20 0.49 0.44 0.32
DBA-Fusion [29] F+I 0.07 0.27 0.10 0.56 0.11 0.13 0.38 0.19 0.23
DEVO [9] E 0.06 0.06 0.71 0.09 0.08 0.06 0.14 0.44 0.21
DVIO F+I 0.05 0.11 0.06 0.26 0.06 0.07 0.38 0.09 0.14
DEIO E+I 0.07 0.09 0.05 0.06 0.04 0.04 0.06 0.08 0.06
  • Unit: MPE(%), 0.06 means the average error would be 0.06m for 100m motion; Aligning 5 seconds [0-5s] of the estimated trajectory with the ground truth;
DAVIS240C [32]
This is the most classical monocular event-based SLAM benchmark, which provides event and image data with a resolution of 240$\times$180 pixels, as well as IMU data and 200HZ ground truth poses. It contains extremely fast 6-Dof motion and scenes with HDR. The estimated and ground-truth trajectories were aligned with a 6-DOF transformation (in SE3), using 5 seconds [0-5s] of the resulting trajectory. We computed the mean position error (MPE) as percentages of the total traveled distance of the ground truth, which are calculated by the publicly available trajectory evaluation tool [41].
The baseline results (except for DPVO, DBA-Fusion, DEVO, and DVIO) are directly taken from the original works, which also employed the same trajectory alignment protocol. As can be seen from Table I, EVI-SAM achieved the best performance among the non-learning methods. However, developed from the feature-based method (PL-EVIO) to the combined direct and feature-based method (EVI-SAM), it’s evident that traditional methods have reached a performance bottleneck, as even the powerful EVI-SAM only offers marginal improvements. In contrast, learning-based methods (such as DEVO) can achieve performance comparable to EVI-SAM using purely event sensors, highlighting the effectiveness and strength of learning-based approaches. While our learning-optimization combined methods exhibits significantly superior performance compared to other learning-based methods (DPVO, DBA-FUsion,DEVO, DVIO). Compared with DEVO, our proposed DEIO decrease the pose tracking error by up to 71%, thanks to the organic combination of learning and traditional optimization method.
Monocular HKU-dataset [10]
TABLE II: Accuracy Comparison of Our DEIO with Other Image/Event-based Baseline in Mono-HKU Dataset [10]
Resolution Methods Modality vicon _hdr1 vicon _hdr2 vicon _hdr3 vicon _hdr4 vicon _darktolight1 vicon _darktolight2 vicon _lighttodark1 vicon _lighttodark2 vicon _dark1 vicon _dark2 Average
DAVIS346 (346*260) ORB-SLAM3 [42] F 0.32 0.75 0.60 0.70 0.75 0.76 0.41 0.58 failed 0.60 0.61
VINS-MONO [30] F+I 0.96 1.60 2.28 1.40 0.51 0.98 0.55 0.55 0.88 0.52 1.02
DPVO [23] F 0.10 0.05 0.12 0.08 0.14 0.08 0.16 0.10 0.06 0.05 0.09
DBA-Fusion [29] F+I 0.32 0.41 failed failed 0.72 0.55 failed 2.65 3.32 failed 1.33
Ultimate-SLAM [3] E+I 1.49 1.28 0.66 1.84 1.33 1.48 1.79 1.32 1.75 1.10 1.40
E+F+I 2.44 1.11 0.83 1.49 1.00 0.79 0.84 1.49 3.45 0.63 1.41
Mono-EIO [10] E+I 0.59 0.74 0.72 0.37 0.81 0.42 0.29 0.79 1.02 0.49 0.62
PL-EIO [4] E+I 0.57 0.54 0.69 0.32 0.66 0.51 0.33 0.53 0.35 0.38 0.49
PL-EVIO [4] F+E+I 0.17 0.12 0.19 0.11 0.14 0.12 0.13 0.16 0.43 0.47 0.20
DEVO [9] E 0.11 0.07 0.12 0.07 0.97 0.12 0.15 0.12 0.07 0.07 0.19
DVIO F+I 0.10 0.18 0.31 0.17 0.15 0.14 0.14 0.26 0.07 0.35 0.19
DEIO E+I 0.14 0.09 0.16 0.07 0.11 0.10 0.11 0.13 0.05 0.08 0.10
  • Unit: MPE(%); Aligning 5 seconds [0-5s] of the estimated trajectory with the ground truth;
(a) boxes_6dof
(b) vicon_darktolight1
(c) HKU_aggressive_small_flip
Figure 3: Comparison of estimated trajectories on challenge sequences from [32, 10, 16].
This dataset was collected using DAVIS346 (346 $\times$ 260, event, image, and IMU sensor), along with a motion capture system (VICON) to obtain pose ground truth. To mitigate the significant impact of the active infrared (IR) emitters from the VICON cameras on the event camera, an IR filter lens was used to eliminate IR interference. As a result, this dataset exhibits more serious thermal noise than others. Furthermore, all sequences were recorded in HDR scenarios under very low illumination or strong illumination changes by switching the strobe flash on and off. Table II demonstrates that DEIO outperform all the event-based methods and decrease the average pose tracking error by at least 47%. Fig. 3(b) illustrates the estimated trajectory of our DEIO, which shows less drift compared to DEVO, verifying the contribution of IMU integration to maintain low-drifting, metric-scale pose estimation.
Stereo HKU-dataset [16]
TABLE III: Accuracy Comparison of Our DEIO with Other Image/Event-based Baseline in Stereo-HKU Dataset [16]
Methods Modality agg_translation agg_rotation agg_flip agg_walk hdr_circle hdr_slow hdr_tran_rota hdr_agg dark_normal Average
ORB-SLAM3 [42] Stereo F+I 0.15 0.35 0.36 failed 0.17 0.16 0.30 0.29 failed 0.25
VINS-Fusion [43] Stereo F+I 0.11 1.34 1.16 failed 5.03 0.13 0.11 1.21 0.86 1.24
ESVO [2] Stereo E failed failed failed failed failed failed failed failed failed
EnVIO [44] Stereo F+I failed 2.12 2.94 3.38 0.85 0.43 0.37 0.50 failed 1.51
DPVO [23] F 0.07 0.04 0.99 1.17 0.31 0.23 0.67 0.29 failed 0.47
DBA-Fusion [29] F+I 0.13 0.16 0.83 0.37 0.18 failed failed 0.10 0.27 0.29
PL-EVIO [4] E+F+I 0.07 0.23 0.39 0.42 0.14 0.13 0.10 0.14 1.35 0.33
EVI-SAM [17] E+F+I 0.17 0.24 0.32 0.26 0.13 0.11 0.11 0.10 0.85 0.25
ESIO [16] Stereo E+I 0.55 0.78 3.17 1.30 0.46 0.31 0.91 1.41 0.35 1.03
ESVIO [16] Stereo E+F+I 0.10 0.17 0.36 0.31 0.16 0.11 0.10 0.10 0.42 0.20
EVO [1] E failed failed failed failed failed failed failed failed failed
DEVO [9] E 0.06 0.05 0.71 0.90 0.39 0.08 0.08 0.26 0.06 0.29
DVIO F+I 0.05 0.06 0.19 1.33 0.15 0.08 0.07 0.04 0.12 0.23
DEIO E+I 0.06 0.09 0.20 0.48 0.14 0.07 0.09 0.06 0.11 0.15
  • Unit: MPE(%) ; Aligning the whole ground truth trajectory with estimated poses; The notations E, F, and I stand for the use of event, frame, and IMU, respectively. The Baseline numbers (DPVO, DEVO) are taken from [9].
This dataset contains stereo event data at 60Hz and stereo image frames at 30Hz with resolution in $346\times 260$, as well as IMU data at 1000Hz. It consists of handheld sequences including rapid motion and HDR scenarios. In Table III, our method outperforms all previous works in terms of average positioning error. Moreover, DEIO beats the DEVO and increase the average accuracy of the sequences up to 48%, especially on agg_walk and agg_flip. DEVO encounters significant errors due to the limitations of relying solely on monocular event camera. As shown in Fig. 3(c), our trajectory aligns perfectly with the ground truth, while DEVO exhibits serious scale loss.
VECtor [45]
TABLE IV: Accuracy Comparison of Our DEIO with Other Image/Event-based Baseline in Vector [45]
Methods Modality corner- slow desk- normal sofa- fast mountain- fast corridors- dolly corridors- walk units- dolly units- scooter average
ORB-SLAM3 [42] Stereo F+I 1.49 0.46 0.21 2.11 1.03 1.32 7.64 6.22 2.81
VINS-Fusion [43] Stereo F+I 1.61 0.47 0.57 failed 1.88 0.50 4.39 4.92 2.05
DPVO [23] F 0.30 0.09 0.07 0.11 0.56 0.54 1.52 1.67 0.61
DBA-Fusion [29] F+I 1.72 0.48 0.43 failed 1.37 0.59 1.23 0.48 0.90
EVO [1] E 4.33 failed failed failed failed failed failed failed 4.33
ESVO [2] Stereo E 4.83 failed failed failed failed failed failed failed 4.83
Ultimate-SLAM [3] E+F+I 4.83 2.24 2.54 4.13 failed failed failed failed 3.44
PL-EVIO [4] E+F+I 2.10 3.66 0.17 0.13 1.58 0.92 5.84 5.00 2.92
ESVIO [16] Stereo E+F+I 1.49 0.61 0.17 0.16 1.13 0.43 3.43 2.85 1.41
EVI-SAM [17] E+F+I 2.50 1.45 0.98 0.38 1.58 1.27 0.59 0.83 1.32
DEVO [9] E 0.59 0.11 0.38 0.37 0.51 1.04 0.48 0.88 0.55
DVIO F+I 0.72 0.26 0.44 0.20 0.54 0.40 1.50 2.37 0.68
DEIO E+I 0.50 0.13 0.44 0.24 0.78 0.74 0.35 0.35 0.44
  • Unit: MPE(%); Aligning the whole ground truth trajectory with estimated poses; The Baseline numbers (DPVO, DEVO) are taken from [9].
This dataset consists of a hardware-synchronized sensor suite that includes stereo event cameras ($640\times 480$), stereo standard cameras ($1224\times 1024$), and IMU. It covers the full spectrum of 6 DoF motion dynamics, environment complexities, and illumination conditions for both small and large-scale scenarios. As seen in Table IV, our proposed DEIO achieves outstanding results on average. It beats all image-based baselines with high-quality frames, and even outperforms Ultimate-SLAM, PL-EVIO, ESVIO, and EVI-SAM on over 75% of the sequences, which utilize event, image and IMU. More Specifically, DEIO also performs better on average than DEVO on large-scale sequences thanks to the complementarity between event and IMU sensor. While the DEVO, monocular visual-only method encounters challenges in these low-texture scenes, and it is also limited by scale ambiguity and drift.
TUM-VIE [33]
TABLE V: Accuracy Comparison of Our DEIO with Other Event-based Baseline in TUM-VIE [33]
Methods Modality mocap- Average
1d-trans 3d-trans 6dof desk desk2
EVO [1] E 7.5 12.5 85.5 54.1 75.2 47.0
ESVO [2] Stereo E 12.3 17.2 13.0 12.4 4.6 11.9
ES-PTAM [46] Stereo E 1.05 8.53 10.25 2.5 7.2 5.9
DH-PTAM [22] Stereo E+F 10.3 0.7 2.4 1.6 1.5 3.9
Ultimate-SLAM [3] E+F+I 3.9 4.7 35.3 19.5 34.1 19.5
DEVO [9] E 0.5 1.1 1.6 1.7 1.0 1.2
DEIO E+I 0.4 1.1 1.4 1.4 0.7 1.0
  • Unit:ATE/ RMSE(cm); Aligning the whole ground truth trajectory with estimated poses; The Baseline numbers (EVO, ESVO, and ES-PTAM) are taken from [46], while DH-PTAM and Ultimate-SLAM are sourced from [22], and DEVO is from [9].
This dataset was recorded with stereo high-resolution event cameras ($1280\times 720$) mounted on a helmet, capturing footage from an egocentric viewpoint. We evaluated our method on sequences recorded in a motion capture room that featured various 6DOF camera movements. The results can be seen in Table V. DEIO outperforms all previous works on four out of five sequences, even though DH-PTAM [22] utilizes four cameras of the setup (stereo events and stereo images).
UZH-FPV [47]
TABLE VI: Accuracy Comparison of Our DEIO with Other Image/Event-based Baseline in UZH-FPV [47]
Methods Modality Indoor_forward Average
3 5 6 7 9 10
VINS-Fusion [43] Stereo F+I 0.84 failed 1.45 0.61 2.87 4.48 2.45
ORB-SLAM3 [42] Stereo F+I 0.55 1.19 failed 0.36 0.77 1.02 0.78
EVO [1] E failed failed failed failed failed failed
DPVO [23] F failed failed failed failed failed failed
VINS-MONO [30] F+I 0.65 1.07 0.25 0.37 0.51 0.92 0.63
DBA-Fusion [29] F+I failed failed failed failed failed failed
Ultimate SLAM [3] E+F+I failed failed failed failed failed failed
PL-EVIO [4] E+F+I 0.38 0.90 0.30 0.55 0.44 1.06 0.60
DEVO [9] E 0.37 0.40 0.31 0.50 0.61 0.52 0.45
DVIO F+I failed failed failed failed failed failed
DEIO E+I 0.39 0.36 0.33 0.32 0.59 0.55 0.42
  • Unit: MPE(%); Aligning the whole ground truth trajectory with estimated poses; The Baseline numbers (DPVO, DEVO) are taken from [9].
This dataset is a high-speed, aggressive dataset collected by monocular DAVIS346, which includes fast laps around a racetrack with drone racing gates, as well as free-form trajectories around obstacles. As shown in Table VI, this dataset is extremely challenging that even the most advanced learning-based VO (DPVO) fails on all sequences. However, our DEIO performed better than the baseline method on average. Indeed, the improvement between DEIO and DEVO is not significant, which can be attributed to the serious IMU bias caused by aggressive motion. Consequently, the gains from the optimization with IMU are restricted. Nevertheless, our DEIO still successfully tracks long stretches of the flight involving to-fro motion, sudden change of direction and loops.
MVSEC [48]
TABLE VII: Accuracy Comparison of Our DEIO with Other Image/Event-based Baseline in MVSEC [48]
Methods Modality Flying_1 Flying_2 Flying_3 Flying_4 Average
ORB-SLAM3 [42] Stereo F+I 5.31 5.65 2.90 6.99 5.21
VINS-Fusion [43] Stereo F+I 1.50 6.98 0.73 3.62 3.21
EVO [1] E 5.09 failed 2.58 failed 3.84
ESVO [2] Stereo E 4.00 3.66 1.71 failed 3.12
Ultimate-SLAM [3] E+F+I failed failed failed 2.77 2.77
PL-EVIO [4] E+F+I 1.35 1.00 0.64 5.31 2.08
ESVIO [16] Stereo E+F+I 0.94 1.00 0.47 5.55 1.99
DPVO [9] F 0.16 0.15 0.08 0.30 0.17
DBA-Fusion [29] F+I 2.20 failed failed failed 2.20
DEVO [9] E 0.26 0.32 0.19 1.08 0.46
DVIO F+I 0.20 0.28 0.16 0.94 0.40
DEIO E+I 0.24 0.21 0.12 0.78 0.34
  • Unit: MPE(%); Aligning the whole ground truth trajectory with estimated poses; The Baseline numbers (DPVO, DEVO) are taken from [9].
This dataset is collected by stereo DAVIS346. Specifically, we selected the indoor_flying sequence, which features arbitrary 6-DOF motion from a drone flying in a motion capture room. The rapid forward and backward movements of the drone, combined with hovering pauses that generate almost no events, make this scenario particularly challenging. Additionally, the few events generated far from the camera is further challenging event-based SLAM. In Table VII, we surpass the event-based baseline, especially, for the Flying_4, we attain a RMSE of 40% lower than DEVO. However, due to the noisy event data of MVSEC, DEIO still performs worse than the learning-based VO (DPVO).
DSEC [49]
TABLE VIII: Accuracy Comparison of Our DEIO with Stereo Event-based Methods in DSEC [49]
Methods Modality dsec_zurich_city_04 Average
a b c d e
ESVO [2] Stereo E 371.1 116.6 1357.1 2676.6 794.9 1032.9
ESVO+IMU [50] Stereo E+I 105.0 66.7 637.9 699.8 130.3 527.9
ES-PTAM [46] Stereo E 131.62 29.02 1184.37 1053.87 75.9 494.9
ESIO [16] Stereo E+I 543.5 295.1 896.2 2977.0 2326.4 1587.8
ESVIO [16] Stereo E+F+I 371.2 445.8 1892.7 921.7 352.0 596.9
DEIO E+I 80.6 35.4 413.8 207.6 86.1 164.5
  • Unit: ATE/RMSE(cm); Aligning the whole ground truth trajectory with estimated poses; Baseline numbers (ESVO and ESVO+IMU) are taken from [50], and ES-PTAM is sourced from [46].
This dataset was collected with stereo event cameras ($640\times 480$) mounted on a car driving through the streets of Switzerland. Since the DSEC dataset lacks ground truth trajectories, we used the odometry trajectories provided by [46] which is obtained through a LiDAR-IMU-based method, as ground truth. As can be seen from the Table VIII, our DEIO beats the event-only methods (ESVO, ESVO+IMU, and ESIO) by large margins on all sequences (at least 66.7% lower RMSE). Fig. 4 presents a qualitative comparison of the estimated trajectories of our DEIO and these stereo event-based baselines. The results from our DEIO align more closely with the ground truth, even though our employ the monocular setup while the others with stereo setup. This proves that the DEIO can obtain comparable scale estimation to these stereo setup and shows better state estimation result. What’s more, as can be seen from the results, the purely event-based odometry (ESVO) exhibits the most significant drift. Although the other stereo event+IMU fusion methods do not perform as well as ours, they maintain minimal drift, emphasizing the importance of complementarity between visual events and IMU sensors.
(a) dsec_zurich_city_04_a
(b) dsec_zurich_city_04_e
(c) dsec_zurich_city_04_b
(d) dsec_zurich_city_04_c
(e) dsec_zurich_city_04_d
Figure 4: Comparison of Estimated trajectories on the DSEC dataset [49].
Figure 5: (a) Our driving platform [8]; (b)The estimated trajectory of our DEIO in the night driving scenarios and its comparison against the GNSS-INS-RTK as ground-truth; The image frame is for visualization-only;
Figure 6: (a) Our quadrotor platform [16]; (b) The estimated trajectory of our DEIO in the quadrotor flight and its comparison against the ground truth; (c) The position, orientation, and the corresponding errors of DEIO in quadrotor flight compared with the ground truth from VICON.
IV-B Test on Night Driving Scenarios
Driving scenarios, especially in nighttime driving environments, pose challenges for event-based state estimation because forward motion typically generates much fewer events events in the center of the image compared to the surrounding area, while the apparent motion at the center of the image is minimal. What’s more, the movement of vehicles, such as sharp turns, sudden stops, and other abrupt movements can further complicate the event-based estimator. In this section, we selected the Dense_street_night_easy_a sequences of the ECMD dataset [8], which feature numerous flashing lights from vehicles, street signs and the surrounding buildings, as well as many moving vehicles, making event-based SLAM more difficult. This dataset were recorded with two pairs of stereo event cameras ($640\times 480$ and $346\times 260$) on a car driven through various road conditions such as streets, highways, road, tunnels in Hong Kong. Our DEIO runs on the event from the DAVIS346 and the IMU sensor, while image frame output of the DAVIS346 is only used for illustration purposes. Fig. 5(b) shows small drift with 4.7m error of our estimated trajectory on the 620m drive. To the best of our knowledge, we present the first results on pose tracking for night driving scenarios using event and IMU odometry. The earliest attempt in this area was ESVIO [16, 8], which utilized a combination of stereo events, images, and IMU data, while the DEIO is monocular setup.
IV-C Test on Dark Quadrotor-Flight
In this section, we evaluate our DEIO in a dark quadrotor flight experiment. The quadrotor is commanded to track a circle pattern with 1.5m in radius and 1.8m in height, shown in Fig. 6(b). The illuminance in the environment is quite low, resulting in minimal visual information captured by the onboard camera (Fig. 6(a)). The total length of the trajectory is 60.7m, with a MPE of 0.15 and an average pose tracking error of 9cm. We further illustrate the estimated trajectories (translation and rotation) of our DEIO against the ground truth, as well as their corresponding errors (Fig. 6(c)). The translation errors in the X, Y, and Z dimensions are all within 0.5m, while the rotation error of the Roll and Pitch dimensions are within 6°, and the one in the Yaw dimension is within 10°. To the best of our knowledge, this is also the first implementation of event-inertial odometry for pose tracking in dark flight environments, while the previous works [3, 16] reply on the image-aided event-inertial estimators.
IV-D Ablation Study
V Conclusion
In this paper, we proposes DEIO which is a Learning-based event-inertial odometry that tightly integrates the deep event-based bundle adjustment with IMU in sliding window graph-based optimization. Evaluation on nine challenging event-based benchmarks demonstrates the superior performance of DEIO compared to image-based and event-based baselines. We show that learning-optimization-combined framework for SLAM is a promising direction. In future work, we might further explore the IMU-bias online learning and loop closure mechanisms for learning-based event-SLAM with global BA to enhance the robustness and efficiency of the system.

### v3

@@
 EVO [1] presents a geometric 2D-3D model alignment, utilizing the 2D image-like event representation and the 3D map reconstructed from EMVS[12].
-It can perform on small-scale scenes, but it is highly parameter-sensitive and requires to run in the scene that is planar to the sensor, up to several seconds, for bootstrapping the system.
 ESVO [2] is the first stereo event-based VO method to estimate the ego motion through the 2D-3D edge registration on time surface.
-However, it barely operates in real-time in low resolution event camera and also faces limitations due to rigorous initialization as well as unreliable pose tracking.
 What’s more, some approaches reply on a pre-build photometric 3D map [13] or restrict the motion type to rotation-only [14].
-However, most of them are prone to failure in challenging scenarios (such as fast motion), which negates the unique advantages of event cameras.
+However, most of them are prone to failure in challenging scenarios (such as fast motion).
 To enhance the robustness of purely event-based VO, existing event-based SLAM methods have demonstrated good performance by incorporating additional sensors.
 Notably, event-inertial integration is a widely used approach to address the limitations of event-only SLAM, which provides scale awareness and continuity of estimation with minimal setup requirements.
-A monocular event-inertial odometry (EIO) is presented in [10], which employs the event-corner features with IMU measurement to deliver real-time and accurate 6-DoF state estimation.
-What’s more, this EIO framework can provide good global consistency through loop closure of the event-based representation.
+Zhu et al.
+[15] proposed the first event-inertial odometry (EIO) which fuses events with IMU through the Extended Kalman Filter.
+A monocular EIO is presented in [10], which employs the event-corner features with IMU measurement to deliver real-time and accurate 6-DoF state estimation.
 Ultimate-SLAM [3] and PL-EVIO [4] investigate the complementary nature of events and images to present an event-image-IMU odometry (EVIO).
@@
 EVI-SAM [18] introduced the first event-based hybrid pose tracking framework, merging the feature-based and direct-based methods.
-Additionally, it consistently delivers excellent mapping results in a variety of challenging scenarios, demonstrating performance comparable to learning-based and NeRF-based mapping methods.
-Despite such remarkable success, current event-based SLAM systems still fall short of the robustness required for certain real-world applications, and their performance has hit a bottleneck, this directs toward the exploration of deep learning approaches.
+Additionally, it consistently delivers excellent mapping results in a variety of challenging scenarios, demonstrating performance comparable to learning-based and NeRF-based dense mapping methods.
+Despite such remarkable progress, current event-based SLAM systems still fall short of the robustness required for certain real-world applications, and their performance has hit a bottleneck, this directs toward the exploration of deep learning approaches.
 II-B Learning-based Approaches for Event-based SLAM
@@
 Klenk, et al.
-[9] extends the DPVO [23] to accommodate the event modality through voxel-based representation, which are able to demonstrate generalization from simulation to several real-world event-based benchmarks.
+[9] extends the DPVO [24] to accommodate the event modality through voxel-based representation, which are able to demonstrate great generalization from simulation to several real-world event-based benchmarks.
 It shows good performance compared to both classical and deep-learning baselines.
-However, their results reveal a significant scale loss/ambiguity, and they requires the careful selection of the input data.
-Pellerito, et al.
-[24] introduces an end-to-end VO which also builds upon DPVO using complicated feature encoders to fuse asynchronous events with image data.
-However, it requires a minimum of 1.2 million events between each pair of frames and discards additional frames if this requirement is not met.
-More importantly, it failed to evaluate their methods under fast motion or HDR scenes, and could only operate in very limited amount of real event data without challenging situations.
-Our framework is closely related to [23, 9] but extends these works by integrating trainable event representations with IMU through our learning-optimization-combined strategy to improve the performance and robustness.
-Additionally, it is important to note that a visual-only (event-only) system, such as DPVO [23] and DEVO [9] has inherent limitations, failures can manifest in numerous ways, such as in low-texture environment, divergence in the optimization algorithm, scale ambiguity, and etc.
-Nevertheless, there is a absence of an implementation that incorporates learning-based event SLAM with IMU fusion framework.
+RAMP-VO [25] introduces an end-to-end VO which also builds upon DPVO using feature encoders to fuse events and image data.
+It achieves robust zero-shot transfer performance to real data despite being trained only on synthetic dataset.
+Our framework is closely related to [24, 9] but extends these works by integrating trainable event-based optical flow with IMU through our learning-optimization-combined strategy to improve the performance and robustness.
+It is important to note that visual-only (or event-only) systems have inherent limitations, with failures manifesting in various ways, such as in low-texture environments, optimization algorithm divergence, and scale ambiguity.
+Visual-IMU integration is the most common solution to address these limitations, which providing scale awareness and continuous estimation with minimal setup.
+Nevertheless, there is a absence of implementation that combines learning-based event SLAM with IMU.
 III Methodology